# Chinese Academy of Sciences, Beijing 100190, China

### v2

@@ 2.2.2 Kinetic equation to $\mathcal{O}(\epsilon^{1})$
where the upper (lower) signs and indices refer to $\Phi$ ($\Phi^{\dagger}$). The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon^{1})$. Therefore, the general statistical function $S^{F}(k,x)$ can be expressed as
@@
Expand the kinetic equation for $S^{F}(k,x)$ to $O(\epsilon)$, we obtain
@@
Substituting the general statistical function $S^{F}(k,x)$, including the $O(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $O(\epsilon)$:
$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=O\left(\epsilon^{2}\right),$ (26)
$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=O\left(\epsilon^{2}\right),$ (27)
$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (28)
$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=O\left(\epsilon^{2}\right),$ (29)
@@
$\Delta_{L/R}^{\kappa}$ $=O\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=O\left(\epsilon^{2}\right),$ (30)
@@ 2.2.3 Kinetic equation to $\mathcal{O}(\epsilon^{2})$
At $O(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
@@ 2.3 QKEs for Dirac and Majorana neutrinos
[…]
@@ 3.3 Numerical Results of Neutrino Decoupling
where $\rho_{\nu}$ is the total neutrino energy density, obtained by solving the QKEs, and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is near that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, leading to a slightly increased $N_{\rm eff}$. The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(3)$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
@@
[…]
@@ 4.1 The Weak Correction from Neutrino Spectral Distortions
×$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
@@
[…]
@@
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between the neutron and proton. The first term in (74) corresponds to the $n\to p+e$ processes (66) or (67), depending on whether $E-\Delta$ is positive or negative, while the second term corresponds to the $n+e\to p$ process (68). Similarly, the terms in (75) can be interpreted. Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $-E_{\nu}$ when it is in the final state. The term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ provides the corresponding distribution or Pauli-blocking factor, following the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
@@
×$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,,$ (79)
where $\Theta(x)$ is the Heaviside step function, and $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ represent the spectral distortions of electron neutrinos and antineutrinos, respectively. For a given process, the various sgn and $\Theta$ functions together characterize the nature and state of the neutrino particle, indicating whether it is a neutrino or an antineutrino and whether it is in the initial or final state. Additionally, we note that $T_{\rm CM}$ and $T_{\nu}$ exhibit slight differences after neutrino decoupling. If we ignore this difference and set $\xi_{\nu}=0$, resulting in $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (4.1) coincides with Eq. (29) in Ref. Froustey:2019owm, where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
@@
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Ref. Pitrou:2018cgg, and are not repeated here for brevity. It is evident that the zero-temperature radiative correction modifies only the coefficients of the $\chi_{\pm}$ functions. The terms $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can be derived simply by replacing $\chi_{\pm}$ with $\delta\chi_{\pm}$. In principle, other corrections are also influenced by the spectral distortions of neutrinos and antineutrinos. However, these effects are expected to be negligible, as the corrections themselves are already sub-leading. Thus, we disregard these interferences in this work.
@@ 4.2 Results of BBN Abundances
In this subsection, we present the resulting light element abundances, incorporating all corrections discussed previously. Specifically, we have implemented the corrections induced by neutrino spectral distortions into the public BBN code PRIMAT Pitrou:2018cgg and computed the abundances of the light elements. For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}h^{2}=0.024$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and number density of isotope $i$, respectively, and $n_{b}$ is the total baryon number density. We also customarily denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
@@
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized to the results for $\xi_{\nu}=0$. The ${}^{3}{\rm He}$ notation refers to ${}^{3}{\rm He}+\rm T$, and ${}^{7}{\rm Li}$ denotes ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. All optional weak rate corrections in PRIMAT (see Ref. Pitrou:2018cgg) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details. The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
@@
Figure 5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (80) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of ${}^{3}{\rm He}$ refers to ${}^{3}{\rm He}+\rm T$, while ${}^{7}{\rm Li}$ corresponds to ${}^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. Since the abundances of different elements can vary by several orders of magnitude, we normalize the element abundances to their respective values at $\xi_{\nu}=0$. Consequently, all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We observe that the abundances of $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$, and ${}^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the corresponding reduction in the neutron-to-proton number density ratio. Furthermore, the variation in ${}^{4}{\rm He}$ abundance is most sensitive to changes in $\xi_{\nu}$, while the ${}^{3}{\rm He}$ abundance is the least sensitive. Additionally, Fig. 5 shows the constraints on helium abundance derived from the CMB and EMPRESS experiments:
$Y_{P}|_{\rm CMB}=0.242\pm 0.012,\quad Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (83)
@@
as represented by the light and dark blue shaded regions. Our results indicate that the constraints on $Y_{P}$ within the $68\%$ confidence interval can be translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
@@
${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,.$ (85)
Clearly, the EMPRESS experiment favors a non-zero positive $\xi_{\nu}$, while the CMB experiment remains compatible with the Standard Model case of $\xi_{\nu}=0$.
@@
It is also interesting to discuss the effects of corrections induced by neutrino spectral distortions on BBN. We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ , $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This treatment is the most common in current BBN studies; (3) the actual SD case, which utilizes the actual $N_{\rm eff}$, accounting for the non-instantaneous decoupling effect and incorporating actual density distributions with spectral distortions for neutrinos and antineutrinos. This case corresponds to the results shown in Fig. 5.
Using the BBN results in the instantaneous decoupling limit as a benchmark, we present the relative differences for the no-SD and actual-SD cases, indicated by dotted and solid curves in Fig. 6, respectively. Note that we have plotted the negative of the relative differences for ${}^{7}{\rm Li}$ in Fig. 6 due to the specific nuclear reaction rates utilized in PRIMAT. We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 1$ in table 1, along with the associated relative variations compared to the baseline abundance in table 2. Our results with $\xi_{\nu}=0$ are consistent with those reported in Froustey:2020mcq , which also considered the effects of neutrino spectral distortions on BBN within the standard model (SM). As shown in Fig. 6 and the accompanying tables, the actual-SD and no-SD cases yield nearly identical light element abundances for $\xi_{\nu}=-0.1$. The BBN abundances in the actual-SD case show minimal dependence on $\xi_{\nu}$, while the traditional no-SD treatment predicts a stronger dependence for both elements. Notably, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the instantaneous decoupling limit. In contrast, the more precise results with spectral distortions indicate positive variations in helium abundance, demonstrating that incorporating neutrino spectral distortions is crucial for accurate predictions of BBN abundances.

### v3

@@ 2.2.2 Kinetic equation to $\mathcal{O}(\epsilon)$
where the upper (lower) signs and indices refer to $\Phi$ ($\Phi^{\dagger}$). The dispersion relations for the functions $F_{L,R}(k,x)$ and $\Phi(k,x)$, as well as for the other spinor components of $S^{F}(k,x)$, are modified by neutrino masses and interactions at $\mathcal{O}(\epsilon)$. Therefore, the general statistical function $S^{F}(k,x)$ can be expressed as
@@
Expand the kinetic equation for $S^{F}(k,x)$ to $\mathcal{O}(\epsilon)$, we obtain
@@
Substituting the general statistical function $S^{F}(k,x)$, including the $\mathcal{O}(\epsilon)$ corrections, into the kinetic equation and decomposing into scalar, vector, and tensor components, we derive the dispersion relations for $F_{L,R}(k,x)$ and $\Phi(k,x)$ at $\mathcal{O}(\epsilon)$:
$\left(k\cdot\hat{\kappa}\right)F_{R}-\frac{1}{2}\left\{\Sigma_{L}^{\kappa},F_{R}\right\}$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (27)
$\left(k\cdot\hat{\kappa}\right)F_{L}-\frac{1}{2}\left\{\Sigma_{R}^{\kappa},F_{L}\right\}$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (28)
$\left(k\cdot\hat{\kappa}\right)P_{+}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{R}^{\kappa}P_{+}^{ij}F_{T}^{j}+P_{+}^{ij}F_{T}^{j}\Sigma_{L}^{\kappa}\right)$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (29)
$\left(k\cdot\hat{\kappa}\right)P_{-}^{ij}F_{T}^{j}-\frac{1}{2}\left(\Sigma_{L}^{\kappa}P_{-}^{ij}F_{T}^{j}+P_{-}^{ij}F_{T}^{j}\Sigma_{R}^{\kappa}\right)$ $=\mathcal{O}\left(\epsilon^{2}\right),$ (30)
@@
$\Delta_{L/R}^{\kappa}$ $=\mathcal{O}\left(\epsilon^{2}\right)\ \ \ \ \,,\ \ \ \ \ \Delta_{T}^{i}=\mathcal{O}\left(\epsilon^{2}\right),$ (31)
@@ 2.2.3 Kinetic equation to $\mathcal{O}(\epsilon^{2})$
At $\mathcal{O}(\epsilon^{2})$, the kinetic equation $S^{F}(k,x)$ becomes
@@ 2.3 QKEs for Dirac and Majorana neutrinos
$\Pi^{\pm}(\vec{k})$ $=\int_{0}^{\infty}dk^{0}\,\hat{\Pi}^{\pm}(k^{0},\vec{k})\,\delta(k^{0}-\left|\vec{k}\right|)\,$, (48)
$\bar{\Pi}^{\pm}(\vec{k})$ $=-\int_{-\infty}^{0}dk^{0}\,\hat{\Pi}^{\mp}(k^{0},-\vec{k})\,\delta(k^{0}+\left|\vec{k}\right|)\,.$ (49)
@@ 3.3 Numerical Results of Neutrino Decoupling
where $\rho_{\nu}$ is the total neutrino energy density, obtained by solving the QKEs, and $\rho_{\gamma}$ is the photon energy density. For the SM in the instantaneous decoupling limit, $N_{\rm eff}$ equals the number of neutrino flavors. However, since the temperature of neutrino decoupling is near that of electron-positron annihilation, neutrinos are also heated through weak interactions with electrons and positrons, leading to a slightly increased $N_{\rm eff}$. The most precise calculation for $N_{\rm eff}$ in the SM, including neutrino oscillations, matter effects, FTQED corrections up to $\mathcal{O}(e^{3})$, and neutrino-electron and neutrino-neutrino collision terms, gives $N_{\rm eff}^{\rm SM}=3.0440\pm 0.0002$ Froustey:2020mcq ; Bennett:2020zkv .
@@
where the second equality uses the condition $\xi_{\nu_{e}}=\xi_{\nu_{\mu}}=\xi_{\nu_{\tau}}=\xi_{\nu}$. Beyond the instantaneous decoupling limit, we explore the non-instantaneous decoupling corrections, $\delta N_{\rm eff}^{\rm dec}=N_{\rm eff}-N_{\rm eff}^{\rm inst}$, in the presence of the primordial neutrino asymmetry $\xi_{\nu}$. In the left panel of Fig. 2, we show $\delta N_{\rm eff}^{\rm dec}$ as a function of $\xi_{\nu}$. For the SM case with $\xi_{\nu}=0$, our calculation yields $\delta N_{\rm eff}^{\rm dec,SM}=0.0441$, consistent with the literature result, $\delta N_{\rm eff}^{\rm dec,SM}=0.0440\pm 0.0002$. [^2] The $\sim 0.0001$ difference between our result and the central value in previous literature arises mainly from different values of $N_{y}$ used for discretizing the neutrino momentum grid. Our result agrees with previous results using $N_{y}=40$; see Table 2 in Bennett:2020zkv . For non-zero $\left|\xi_{\nu}\right|$, $\delta N_{\rm eff}^{\rm dec}$ increases beyond $\delta N_{\rm eff}^{\rm dec,SM}$, as expected since a higher $\left|\xi_{\nu}\right|$ results in more energy transfer to neutrinos from electrons and positrons. We find that $\delta N_{\rm eff}^{\rm dec}$ as a function of $\xi_{\nu}$ can be well fitted by $\delta N_{\rm eff}^{\rm dec}=\delta N_{\rm eff}^{\rm dec,SM}+0.0102\,\xi_{\nu}^{2}$, as shown in the left panel of Fig. 2. In conclusion, the final $N_{\rm eff}$ after neutrino decoupling, in the presence of primordial neutrino asymmetry $\xi_{\nu}$, is given by
@@ 4.1 The Weak Correction from Neutrino Spectral Distortions
$\times f_{a}(E_{a})f_{b}(-E_{b})f_{e}(\alpha_{e}{E}_{e})\tilde{f}_{\nu}(E_{\nu}),$
@@
To leading order, two additional approximations are typically made. First, we neglect neutrino spectral distortions and use the Fermi-Dirac distribution with the corresponding chemical potentials $f_{\nu}^{\pm}(E)\equiv\frac{1}{e^{E/T_{\nu}\mp\xi_{\nu}}+1}$ for neutrinos and antineutrinos, where $T_{\nu}$ is the effective temperature of the neutrino, defined as the temperature that satisfies $3\frac{7}{8}\frac{\pi^{2}}{15}T_{\nu}^{4}=\rho_{\nu}$. Furthermore, since the temperature during the BBN epoch (approximately $0.1\,{\rm MeV}$) is much smaller than the masses of the neutron and proton (approximately $1\,{\rm GeV}$), we can work within the infinite nucleon mass limit, also known as the Born approximation. Under these two approximations and following the notation of Pitrou:2018cgg , the reaction rates are significantly simplified as follows:
@@
where $\Delta\equiv m_{n}-m_{p}$ is the energy gap between the neutron and proton. The first term in (75) corresponds to the $n\to p+e$ processes (67) or (68), depending on whether $E-\Delta$ is positive or negative, while the second term corresponds to the $n+e\to p$ process (69). Similarly, the terms in (76) can be interpreted. Here, $E^{\mp}_{\nu}$ is equal to the energy of the neutrino $E_{\nu}$ when the neutrino is in the initial state, and $E_{\nu}$ when it is in the final state. The term $f_{\nu}^{\pm}(E^{\mp}_{\nu})$ in the function $\chi^{\pm}$ provides the corresponding distribution or Pauli-blocking factor, following the relation $1-f_{\nu}^{+}(E_{\nu})=f_{\nu}^{-}(-E_{\nu})$.
@@
$\times\left(\frac{1+\Theta(\pm E^{\mp}_{\nu})\delta\varrho_{\nu}+\Theta(\mp E^{\mp}_{\nu})\delta\varrho_{\bar{\nu}}}{e^{\left|E^{\mp}_{\nu}\right|/T_{\rm CM}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}-\frac{1}{e^{\left|E^{\mp}_{\nu}\right|/T_{\nu}\mp{\rm sgn}(E^{\mp}_{\nu})\xi_{\nu}}+1}\right)\,$, (80)
where $\Theta(x)$ is the Heaviside step function, and $\delta\varrho_{\nu}$ and $\delta\varrho_{\bar{\nu}}$ represent the spectral distortions of electron neutrinos and antineutrinos, respectively. For a given process, the various sgn and $\Theta$ functions together characterize the nature and state of the neutrino particle, indicating whether it is a neutrino or an antineutrino and whether it is in the initial or final state. Additionally, we note that $T_{\rm CM}$ and $T_{\nu}$ exhibit slight differences after neutrino decoupling. If we ignore this difference and set $\xi_{\nu}=0$, resulting in $\delta\varrho_{\nu}=\delta\varrho_{\bar{\nu}}$, then Eq. (4.1) coincides with Eq. (29) in Froustey:2019owm , where corrections to the weak rates from neutrino spectral distortions in the SM were considered.
@@
The explicit expressions for the Fermi function $F_{\pm}$ and the radiative correction factor $R(E,k_{\rm max})$ can be found in, e.g., Pitrou:2018cgg , and are not repeated here for brevity. It is evident that the zero-temperature radiative correction modifies only the coefficients of the $\chi_{\pm}$ functions. The terms $\delta\Gamma^{\rm RC0,SD}_{n\to p}$ and $\delta\Gamma^{\rm RC0,SD}_{p\to n}$ can be derived simply by replacing $\chi_{\pm}$ with $\delta\chi_{\pm}$. In principle, other corrections are also influenced by the spectral distortions of neutrinos and antineutrinos. However, these effects are expected to be negligible, as the corrections themselves are already sub-leading. Thus, we disregard these interferences in this work.
@@ 4.2 Results of BBN Abundances
In this subsection, we present the resulting light element abundances, incorporating all corrections discussed previously. Specifically, we have implemented the corrections induced by neutrino spectral distortions into the public BBN code PRIMAT Pitrou:2018cgg and computed the abundances of the light elements. For this computation, we utilized the latest values of relevant physical and cosmological parameters; for instance, the neutron lifetime is taken to be $\tau_{n}=879.4\,\text{s}$, the axial coupling of the nucleons is $g_{A}=1.2756$, and the baryon density is $\omega_{b}=0.0224$. For convenience, we define the mass fraction of isotope $i$ as $X_{i}\equiv A_{i}\frac{n_{i}}{n_{b}}$, where $A_{i}$ and $n_{i}$ are the baryon number and number density of isotope $i$, respectively, and $n_{b}$ is the total baryon number density. We also customarily denote $Y_{P}\equiv X_{{}^{4}{\rm He}}$ and $i/\rm H\equiv n_{i}/n_{{}^{1}\rm H}$.
@@
Figure 5: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$, normalized with the results for $\xi_{\nu}=0$. The baryon density is fixed at $\omega_{b}=0.0224$, and the nuclear rates of the BBN are driven by the PRIMAT approach. The $^{3}{\rm He}$ notation refers to $^{3}{\rm He}+\rm T$, and $^{7}{\rm Li}$ denotes $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. All optional weak rate corrections in PRIMAT (see Pitrou:2018cgg ) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details. The $68\%$ constraints on the helium abundance $Y_{P}$ from the CMB ($Y_{P}=0.242\pm 0.012$), the PDG recommended value ($Y_{P}=0.245\pm 0.003$) and the recent EMPRESS experiment ($Y_{P}=0.2370^{+0.0034}_{-0.0033}$) are indicated by the light and dark blue shaded regions, respectively.
@@
Fig. 5 illustrates the resulting light element abundances, incorporating all weak rate corrections from (81) and the QED corrections up to $\mathcal{O}(e^{3})$ for plasma thermodynamics and certain nuclear rates, as a function of the neutrino degeneracy parameter $\xi_{\nu}$. The abundance of $^{3}{\rm He}$ refers to $^{3}{\rm He}+\rm T$, while $^{7}{\rm Li}$ corresponds to $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. Since the abundances of different elements can vary by several orders of magnitude, we normalize the element abundances to their respective values at $\xi_{\nu}=0$. Consequently, all curves in Fig. 5 intersect at $\xi_{\nu}=0$. We observe that the abundances of $\rm D,\,^{3}{\rm He},\,^{4}{\rm He}$, and $^{7}{\rm Li}$ decrease with increasing $\xi_{\nu}$ due to the corresponding reduction in the neutron-to-proton number density ratio. Furthermore, the variation in $^{4}{\rm He}$ abundance is most sensitive to changes in $\xi_{\nu}$, while the $^{3}{\rm He}$ abundance is the least sensitive. Additionally, Fig. 5 shows the constraints on helium abundance derived from the CMB observation, the EMPRESS observation and the PDG recommended value:
$Y_{P}|_{\rm CMB}=0.242\pm 0.012,$ (84)
$Y_{P}|_{\rm EMPRESS}=0.2370^{+0.0034}_{-0.0033},$ (85)
$Y_{P}|_{\rm PDG}=0.245\pm 0.003$ (86)
@@
as represented by the blue shaded regions. Through our numerical results with the fixed baryon density $\omega_{b}=0.0224$, the constraints on $Y_{P}$ within the $68\%$ confidence interval can be easily translated into constraints on the neutrino degeneracy parameter $\xi_{\nu}$:
@@
${\rm EMPRESS}:\quad 0.027\leq$ $\,\xi_{\nu}\leq 0.054\,$, (88)
${\rm PDG}:\quad-0.003\leq$ $\,\xi_{\nu}\leq 0.021\,.$ (89)
Clearly, while the CMB observation yields a significantly weaker constraint on the helium abundance $Y_{P}$, the constraint from the EMPRESS observation has similar uncertainties but a much smaller central value with respect to the PDG recommended observation. Consequently, the EMPRESS observation favours a non-zero positive $\xi_{\nu}$, while the CMB and PDG recommended observations remain compatible with the Standard Model case of $\xi_{\nu}=0$.
Key Nuclear Reactions
[…]
3He+D $\to$ 4He+p 3He+4He $\to$ $\gamma 7Be$+n $\to$ 7Li+p 7Li+p $\to$ 4He+4He
Table 1: The key nuclear reactions that possessing different reaction rates in the PRIMAT and NACRE II approaches.
Figure 6: Light element abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$ that driven by the PRIMAT approach (solid curves) and the NACRE II approach (dash-dotted curves), normalized with the results for $\xi_{\nu}=0$ in the PRIMAT approach. The $^{3}{\rm He}$ notation refers to $^{3}{\rm He}+\rm T$, and $^{7}{\rm Li}$ denotes $^{7}{\rm Li}+\rm^{7}Be$, accounting for slow radioactive decays. All optional weak rate corrections in PRIMAT (see Pitrou:2018cgg ) and the new corrections from neutrino spectral distortions are included in the calculations, see text for more details. The $68\%$ constraints on the deuterium abundance $D/H$ of the PDG recommended value ($D/H\times 10^{5}=2.547\pm 0.025$) are indicated by the light red shaded region.
It is also interesting to consider the deuterium abundance constraint. For this purpose, we first briefly discuss the two most widely used databases for thermonuclear rates relevant to BBN, as these are very relevant to the deuterium abundance predicted by BBN. The first database is tabulated by the public BBN code PRIMAT with an extensive catalogue of more than 400 reactionsPitrou:2018cgg , implemented according to the refined statistical analyses of Refs. Descouvemont:2004cw ; Iliadis:2016vkw ; InestaGomez:2017eya . We will refer to the predicted BBN abundances obtained with this database as the PRIMAT approach, and use it to obtain most of the numerical results in this paper. On the other hand, the NACRE II database Xu:2013fha gives an evaluation of 12 key nuclear reaction rates with mass number $A<16$, based on the nuclear cross sections described by the so-called potential model Angulo:1999zz , as shown in Tab. 1. With the same nuclear rates of the other reactions with the PRIMAT approach, we will denote it as the NACRE II approach.
We present the resulting BBN abundances as a function of the neutrino degeneracy parameter $\xi_{\nu}$ in these two approaches in Fig. 6, where the solid curves refer to the PRIMAT approach, which is identical to the results in Fig. 5, and the dashed curves refer to the NACRE II approach. All curves are normalized using the resulting for $\xi_{\nu}=0$ in the PRIMAT approach, and the $68\%$ constraints on the deuterium abundance $D/H$ of the PDG recommended value ($D/H\times 10^{5}=2.547\pm 0.025$) are indicated by the light red shaded region. Comparing the light element abundances obtained by the two approaches, we find that they give exactly the same predictions for $Y_{P}$ and similar predictions for $^{3}{\rm He}/H$. However, the NACRE II approach predicts a a significantly larger $D/H$ and smaller $^{7}{\rm Li}/H$ than the PRIMAT approach. Consequently, the constraints on the neutrino degeneracy parameter $\xi_{\nu}$ from the PDG constraint $D/H\times 10^{5}=2.547\pm 0.025$ with the fixed baryon density $\omega_{b}=0.0224$ are also significantly different in the two approaches:
${\rm PRIMAT}:\quad-0.088\leq$ $\,\xi_{\nu}\leq-0.057\,$, (90)
${\rm NACRE\,II}:\quad-0.042\leq$ $\,\xi_{\nu}\leq-0.009\,.$ (91)
We see that, while the PRIMAT approach points to a much smaller $\xi_{\nu}$, both approaches favor a negative negative neutrino asymmetry $\xi_{\nu}$. Such results indicate a mild tension between current observations of helium and deuterium abundances. In addition, we note that despite the mild tension between $Y_{P}$ and $D/H$ observations, the preference for positive $\xi_{\nu}$ from the EMPRESS observation on $Y_{P}$ is more reliable, since $Y_{P}$ is more sensitive to the neutrino asymmetry $\xi_{\nu}$, while $D/H$ is more sensitive to the baryon density $\omega_{b}$, which is fixed in our upper analysis. Therefore, the preference for a negative $\xi_{\nu}$ from the $D/H$ observation can be greatly mitigated with a slightly different $\omega_{b}$. Consequently, an overall MCMC analysis combining helium and deuterium observations still indicates a preference for positive $\xi_{\nu}$, as we will see in Sec. 6. For the sake of simplicity, we will also only consider the PRIMAT approach in the following, and leave a detailed comparison between different nuclear reaction rates to our subsequent companion paper Li:2025rjr .
@@
It is also interesting to discuss the effects of corrections induced by neutrino spectral distortions on BBN. We examine the impact of different treatments of the neutrino decoupling process, considering three implementations: (1) the neutrino instantaneous decoupling limit, $i.e.,$ $N_{\rm eff}^{\rm inst}$, as discussed in the previous section, where thermal Fermi-Dirac distributions for neutrinos and antineutrinos are used in the BBN calculations; (2) the no-SD case, which employs the actual $N_{\rm eff}$ and neutrino temperatures, incorporating the non-instantaneous decoupling effect while only using thermal Fermi-Dirac distributions for neutrinos and antineutrinos. This treatment is the most common in current BBN studies; (3) the actual SD case, which utilizes the actual $N_{\rm eff}$, accounting for the non-instantaneous decoupling effect and incorporating actual density distributions with spectral distortions for neutrinos and antineutrinos. This case corresponds to the results shown in Fig. 5.
Using the BBN results in the instantaneous decoupling limit as a benchmark, we present the relative differences for the no-SD and actual-SD cases, indicated by dotted and solid curves in Fig. 7, respectively. Note that we have plotted the negative of the relative differences for $^{7}{\rm Li}$ in Fig. 7 due to the specific nuclear reaction rates utilized in PRIMAT. We report the explicit abundances of helium-4, deuterium, helium-3, and lithium-7 for the degeneracy parameters $\xi_{\nu}=0,\pm 0.1$ in table 2, along with the associated relative variations compared to the baseline abundance in table 3. Our results with $\xi_{\nu}=0$ are consistent with those reported in Froustey:2020mcq , which also considered the effects of neutrino spectral distortions on BBN within the standard model (SM). As shown in Fig. 7 and the accompanying tables, the actual-SD and no-SD cases yield nearly identical light element abundances for $\xi_{\nu}=-0.1$. The BBN abundances in the actual-SD case show minimal dependence on $\xi_{\nu}$, while the traditional no-SD treatment predicts a stronger dependence for both elements. Notably, the no-SD case predicts negative variations for the helium abundance with positive $\xi_{\nu}$ compared to the instantaneous decoupling limit. In contrast, the more precise results with spectral distortions indicate positive variations in helium abundance, demonstrating that incorporating neutrino spectral distortions is crucial for accurate predictions of BBN abundances.